OpenStudy (anonymous):

Anyone good with finding limits? Lim(x->0) x^2cos(1/x). Help is appreciated

myininaya (myininaya):

L'hospital?

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{\cos(\frac{1}{x})}{\frac{1}{x^2}}\]

myininaya (myininaya):

Find derivative of the top Find derivative of the bottom

OpenStudy (anonymous):

doesnt that result in 1 regardless?

OpenStudy (anonymous):

can you walk me through the derivative formula?

myininaya (myininaya):

\[(\cos(\frac{1}{x}))'\] Here you need to know the chain rule.

myininaya (myininaya):

\[(\frac{1}{x})'=(x^{-1})' \] Here you to know the power rule.

myininaya (myininaya):

\[(\frac{1}{x^2})'=(x^{-2})'\] Here you need to also know the power rule.

myininaya (myininaya):

Chain rule goes like this: \[(f(g(x))'=g'(x) \cdot f'(g(x))\] Power rule goes like this: \[(x^n)'=nx^{n-1} \]

OpenStudy (anonymous):

so -2x^-2-1 so -2x^(-3) for the bottom?

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{(\cos(\frac{1}{x}))'}{\frac{-2}{x^3}}\] now we still need to take care of the top

myininaya (myininaya):

\[(f(g(x))'=g'(x) \cdot f'(g(x))\] \[(\cos(\frac{1}{x}))'=(\frac{1}{x})' \cdot (\cos)'(\frac{1}{x})\]

myininaya (myininaya):

So what is the derivative of 1/x?

OpenStudy (anonymous):

x^-1

myininaya (myininaya):

yep what is the derivative of that

OpenStudy (anonymous):

-1/x^2 dx = y' ?

myininaya (myininaya):

right so we have \[(f(g(x))'=g'(x) \cdot f'(g(x)) \] \[(\cos(\frac{1}{x}))'=\frac{-1}{x^2} \cdot (\cos)'(\frac{1}{x}) \]

myininaya (myininaya):

so we still need the derivative of cos

OpenStudy (anonymous):

sin?

myininaya (myininaya):

-sin

myininaya (myininaya):

so we have \[\lim_{x \rightarrow 0}\frac{\frac{1}{x^2}\sin(\frac{1}{x})}{\frac{-2}{x^3}} \]

OpenStudy (anonymous):

is it -2/x^3 or -2/x^-3?

OpenStudy (anonymous):

i guess i made a mistake with the -3, so based off that, is that the final solution then?

myininaya (myininaya):

\[\frac{-1}{2} \lim_{x \rightarrow 0}\frac{ \frac{1}{x^2} \sin(\frac{1}{x})}{\frac{1}{x^2} \cdot \frac{1}{x}}\]

OpenStudy (anonymous):

I'm totally lost on how that is simplified, arent you left with sin(1/x)/1/x, so -1/2* sin(0)?

myininaya (myininaya):

1/x^2 cancels out with the 1/x^2 on bottom is how to simplify

OpenStudy (anonymous):

but that leaves you with division by 0?

OpenStudy (anonymous):

sin(1/x)/1/x= sin(1/0)/1/0

myininaya (myininaya):

but \[\lim_{u \rightarrow \infty }\frac{\sin(u)}{u}=0\]

myininaya (myininaya):

By squeeze thm.

OpenStudy (anonymous):

so my initial answer of 0 was right even though i did it wrong lol?

myininaya (myininaya):

yes -1/2*zero will be zero.

OpenStudy (anonymous):

thank you i will print this out and review all the work, appreciate it

myininaya (myininaya):

http://www.math.washington.edu/~conroy/general/sin1overx/ Here is a cute little visual showing you that x*sin(1/x) is between -x and x. since -x and x both approach 0 as x approaches 0 then x*sin(1/x) approaches 0 as x approaches 0.

myininaya (myininaya):

Someone may started out differently than I did here, but I just remember the rule that as u approaches infinity then sin(u)/u approaches 0.